# S2AM3D: Scale-controllable Part Segmentation of 3D Point Cloud

### v1

@@ Abstract
Part-level point cloud segmentation has recently attracted significant attention in 3D computer vision. Nevertheless, existing research is constrained by two major challenges: native 3D models lack generalization due to data scarcity, while introducing 2D pre-trained knowledge often leads to inconsistent segmentation results across different views. To address these challenges, we propose S2AM3D, which incorporates 2D segmentation priors with 3D consistent supervision. We design a point-consistent part encoder that aggregates multi-view 2D features through native 3D contrastive learning, producing globally consistent point features. A scale-aware prompt decoder is then proposed to enable real-time adjustment of segmentation granularity via continuous scale signals. Simultaneously, we introduce a large-scale, high-quality part-level point cloud dataset with more than 100k samples, providing ample supervision signals for model training. Extensive experiments demonstrate that S2AM3D achieves leading performance across multiple evaluation settings, exhibiting exceptional robustness and controllability when handling complex structures and parts with significant size variations.
@@ 1 Introduction
Part-level point cloud segmentation plays a pivotal role in bridging fine-grained geometric details with high-level semantic understanding, supporting significant applications in 3D content creation, robotic manipulation, and reverse engineering, etc.. Unlike instance detection, which only provides holistic semantics, part-level segmentation enables flexible granularity adjustment and seamless switching between local and global regions. This capability directly influences the feasibility and efficiency of downstream tasks, including part generation, replacement, assembly, and parametric editing. However, high-quality 3D part-level data remains scarce. Compared with the 2D domain, the cost of point cloud annotation is more prohibitive; consequently, existing datasets suffer from limited scale and category diversity, which severely hinders the generalization performance of 3D models [qi2017pointnet, qi2017pointnet++, wang2018sgpn, li2018pointcnn, ma2025p3].
@@
Recent works [liu2023partslip, yang2024sampart3d, liu2025partfield, zhou2024point, zhou2023partslip++, garosi20253d, Sun_2024_CVPR] leverage 2D pre-trained knowledge to enhance the generalization of 3D models, i.e., adopting 2D segmentation on 3D renderings. Segmented multi-view contents are then aggregated via mechanisms like multi-view lifting [garosi20253d, Sun_2024_CVPR, liu2023partslip, zhou2023partslip++] or distillation [yang2024sampart3d, liu2025partfield, zhou2024point] to obtain the final 3D segmentation results. However, cross-view inconsistencies caused by occlusions, slender structures, and complex topology, can lead to accumulated errors that compromise global 3D consistency. To alleviate the inconsistency, some methods [liu2025partfield, zhou2024point] further incorporate 2D priors and provide supervision in the 3D space. Although native 3D data is used to enhance multi-view consistency, these methods still rely too heavily on 2D segmentation results. For example, PartField [liu2025partfield] primarily leverages part proposals from SAM [kirillov2023segment], supplemented with limited 3D supervision to learn multi-view dependencies. Consequently, when 2D segmentation is affected by occlusions and other spatial factors, 3D representations struggle to correct these issues (see Figure 1).
@@
Extensive experiments are conducted to demonstrate the effectiveness of S2AM3D on part-level point cloud segmentation. As shown in Figure 1, our 2D-3D hybrid training recipe allows S2AM3D to achieve comparable performance to P3-SAM [ma2025p3] with far less training data, and significantly outperforms SAMPart3D and PartField. Benefitting from the proposed scale-aware prompt decoder, S2AM3D exhibits remarkable flexibility by naturally extending its capability to point-prompted segmentation, making it a unified 3D part segmentation framework.
@@ 2 Related Work
Native 3D Part Segmentation. Early point cloud/mesh segmentation methods were typically trained under closed-label settings on ShapeNet-Part and PartNet, leveraging pointwise or neighborhood aggregation networks such as PointNet/PointNet++ and the Point Transformer family [qi2017pointnet, qi2017pointnet++, zhao2021point, yi2016scalable, mo2019partnet]. While strong within restricted categories and fixed granularities, these models often struggle under open-domain shapes, long-tailed parts, and cross-source shifts: (i) limited category/hierarchy coverage hampers generalization to unseen objects and rare parts; (ii) geometric-detail and noise discrepancies across sources obscure unified supervision; and (iii) the absence of explicit hierarchical consistency and granularity control constrains editing and interactive scenarios [chang2015shapenet, mo2019partnet, collins2022abo]. These limitations motivate scalable part understanding beyond closed-world assumptions.
@@
2D Transferred Part Segmentation. To mitigate the scarcity of native 3D annotations, many works transfer 2D foundation priors (e.g., SAM, CLIP/GLIP) to 3D [kirillov2023segment, radford2021learning, li2022grounded]. Two mainstream paradigms emerge: (i) multi-view lifting renders shapes to images, applies 2D segmentation or vision–language priors per view, and fuses/back-projects/distills view-level cues into 3D supervision [kirillov2023segment, yang2023sam3d, xu2023sampro3d, liu2023partslip]; and (ii) 2D-to-3D distillation directly trains native 3D networks from image-domain pseudo labels [huang2024segment3d, liu2025partfield, yang2024sampart3d]. Despite reduced labeling cost and open-vocabulary semantics, both paradigms are prone to cross-view inconsistencies on thin/occluded structures, boundary ambiguity that undermines global 3D coherence, and costly post-processing pipelines; moreover, surface-only coverage limits supervision of internal structures [yang2023sam3d, xu2023sampro3d, liu2023partslip, huang2024segment3d]. These observations point to the value of native 3D signals for globally consistent part supervision.
@@
Granularity Controllable Part Segmentation. Existing methods still face significant limitations in achieving flexible control over segmentation granularity. Feature clustering-based approaches (e.g., PartField [liu2025partfield] and SAMPart3D [yang2024sampart3d]) typically rely on post-processing clustering to determine segmentation granularity, resulting in non-continuous and non-intuitive control that hinders real-time fine-grained adjustment. Point prompt-based methods (e.g., Point-SAM [zhou2024point] and P3-SAM [ma2025p3]), while supporting interactive point prompting, lack explicit granularity control mechanisms and cannot precisely regulate segmentation scope through continuous parameters. Furthermore, the segmentation granularity in both types of methods remains fundamentally constrained by their reliance on 2D priors, making it challenging to ensure global consistency in 3D space. These limitations motivate us to explore a new solution that introduces an explicit continuous scale signal coupled with native 3D supervision, ultimately achieving truly reliable and controllable 3D segmentation granularity.
@@ 3.2 Point-Consistent Part Encoder
To aggregate point cloud features with general segmentation priors, the common practice is to train a point cloud encoder by distilling the pre-training knowledge from 2D segmentation models like SAM [kirillov2023segment]. Following PartField  [liu2025partfield], we exploit a voxel-based encoder PVCNN [liu2019point] to extract the point latent and then convert it to the tri-plane [chan2022efficient] ($xy$, yzyz, and $zx$) representation with orthogonal projection. The Tri-plane features $\mathbf{T}\!\in\!\mathbb{R}^{3\times D\times H\times W}$ are then aggregated by a series of Transformer blocks, where $D$ is the feature channel. During training, tri-plane features are rendered from random viewpoints to formulate a 2D latent, which can be supervised with the distillation of SAM. However, relying solely on multi-view 2D distillation tends to introduce multi-view inconsistencies, leading to segmentation problems like boundary artifacts and cross-view conflicts.
@@
To obtain point features $\mathbf{F}$ given the 3D coordinate $(x,y,z)$ from the tri-plane representation [chan2022efficient], we unproject the point to the three feature planes of $\mathbf{T}$. The sampled features are summed across planes and stacked over all points as:
@@ 3.3 Scale-Aware Prompt Decoder
where $\{\omega_{k},\phi_{k}\}$ are learnable and $M$ denotes the number of sinusoidal frequency pairs. $\mathbf{e}(s)$ is then fed to a feature-wise linear modulation [perez2018film] (FiLM) in the channel dimension, which modulates the global feature map with a learnable gate $\alpha$:
$=\text{Linear}\!\big(\mathrm{LN}(\mathbf{e}(s))\big),$ (4)
@@
where Linear() is a linear layer applied after Layer Normalization on $\mathbf{e}(s)$, $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{D}$ are channel-wise FiLM parameters [perez2018film], and $\alpha\in\mathbb{R}$ is a learnable scalar gate. To maintain a consistent scale context across semantic levels, we interleave FiLM with Transformer blocks $T$:
@@ 3.4 Training Objective
We adopt a decoupled training scheme: our point-consistent part encoder is trained with the contrastive objective [khosla2020supervised] in Sec. 3.2, while the scale-aware prompt decoder is optimized subsequently with the encoder frozen.
@@ 4.1 Scalable Data Pipeline
The data pipeline consists of three steps: part annotation, quality filtering, and connectivity refinement. Given a 3D object, we adopt surface-area–proportional sampling and assign the corresponding part label to every sampled point, which follows the part-annotation mining strategy of PartCrafter [lin2025partcrafter]. However, such an automatic strategy inevitably introduces erroneous annotation artifacts, which may mislead the model training. To filter out possible wrong labels, we further propose quality filtering and connectivity refinement.
In the quality filtering stage (see Figure 3(b)), we first manually curate a small-scale dataset containing examples of both reasonable and unreasonable part annotations. We then train a binary PointNet [qi2017pointnet] validator on this dataset. Each point cloud is represented by $N$ points, each defined by a 4D vector $(x,y,z,l)$ encompassing normalized coordinates and its part-label index. Once trained, this validator is applied across the entire dataset to automatically screen and filter out samples with unreasonable annotations, ensuring that only high-quality labeled data proceed to subsequent stages.
We then perform connectivity refinement (see Figure 3(c)): artist annotations may group spatially disconnected yet semantically related regions under the same label, which degrades prompt-segmentation precision and hampers isolating local structures. For the points of each label, we compute the axis-aligned bounding-box diagonal length $d$ and set the DBSCAN [ester1996density] radius to $\epsilon=d\times\epsilon_{\text{factor}}$. If multiple spatial clusters are detected within a single label, we split it into distinct new labels. Finally, we filter and retain objects containing 2 to 50 parts to ensure dataset balance and suitability for part segmentation tasks.
@@ 4.2 Dataset Overview
We mainly collect 3D assets from Objaverse [deitke2023objaverse-xl, deitke2023objaverse], which is one of the largest 3D object datasets. We filter more than 100,000 point cloud instances from this dataset, spanning over 400 categories. This category advantage lies in covering open-world shapes from simple objects to complex machinery, supporting broader geometric and semantic diversity. In total, we annotate around 1.2 million part labels, with an average of 11 parts per object. As shown in Figure 3(a), the part count distribution spans 2-50, exhibiting a long-tailed pattern to ensure balanced coverage of fine- and coarse-grained shapes.
@@ 5.1 Implementation Details
Encoder training. The encoder is initialized with PartField [liu2025partfield] pre-trained parameters and then further optimized with contrastive learning on native point cloud annotations. For each mini-batch, we uniformly sample $\hat{N}=1000$ points to form $\hat{P}$ and compute the contrastive loss on it with temperature $\tau=0.07$. We use AdamW [loshchilov2017decoupled] with a learning rate of $1\times 10^{-5}$. The encoder is trained for 15 epochs, which takes about one day on 8 NVIDIA A6000 GPUs.
Decoder training. In this phase, we freeze the parameters of the encoder and fine-tune the decoder only. We set the layer number of scale modulation $L_{m}$ to 2 and the layer number of bi-directional cross-attention blocks $L_{d}$ to 4. The segmentation loss is a weighted sum of dynamic BCE and Dice, with $\lambda_{\mathrm{bce}}=0.7$ and $\lambda_{\mathrm{dice}}=0.3$. During training, we apply the scale dropout with a probability of 0.1, and at inference, we threshold the probability at $\theta=0.7$. We use AdamW [loshchilov2017decoupled] with a learning rate of $1\times 10^{-4}$. The training batch size is 40. The decoder is trained for 110 epochs with 4 NVIDIA A6000 GPUs.
@@ 5.2 Comparison
Point-SAM [zhou2024point] 31.46 50.23 40.85
P3-SAM [ma2025p3] 35.05 39.98 37.52
Ours 41.70 54.39 48.05
Ours (+scale) 53.75 70.63 62.19
@@
Figure 5: Qualitative comparison of full segmentation (PartObjaverse-Tiny [yang2024sampart3d]). For ease of comparison with our point cloud method, mesh-level outputs are presented as point clouds by uniformly sampling the segmented meshes.
@@
Find3D [ma2025find] 20.76 21.69 21.23
SAMPart3D [yang2024sampart3d] 48.79 56.17 52.48
SAMesh [tang2024segment] - 26.66 -
PartField [liu2025partfield] 51.54 59.10 55.32
P3-SAM [ma2025p3] 58.10 65.39 61.75
Ours 54.72 70.14 62.43
@@
Evaluation Datasets. We choose PartObjaverse-Tiny [yang2024sampart3d] and PartNet-E [liu2023partslip] for experimental evaluation. PartObjaverse-Tiny [yang2024sampart3d] is a compact subset of Objaverse [deitke2023objaverse, deitke2023objaverse-xl], containing 200 samples across 8 categories with manual part annotations. PartNet-E [liu2023partslip] provides 1,906 point cloud shapes across 45 categories, which can be used to assess cross-category generalization.
@@
Baselines. For interactive segmentation, we compare two point-prompt baselines: Point-SAM [zhou2024point] and P3-SAM [ma2025p3]. For fair comparison, all methods are evaluated with their default input configurations: Point-SAM [zhou2024point] employs both coordinates and color information, while P3-SAM [ma2025p3] utilizes coordinates and surface normals. In contrast, our approach requires only XYZ coordinates. For full segmentation, we select current state-of-the-art methods, including Find3D [ma2025find], SAMPart3D [yang2024sampart3d], SAMesh [tang2024segment], PartField [liu2025partfield], and P3-SAM [ma2025p3].
@@
Metrics. We use Intersection over Union (IoU) as the evaluation metric. The metric is computed across all parts of a single object to obtain the per-sample score, which is subsequently averaged over the entire dataset. For interactive segmentation, we follow the strategy in Point-SAM [zhou2024point], which selects the prompt point for each part as the interior point farthest from its boundary. All methods are evaluated on 10,000 points normalized to the unit sphere. For full segmentation, we follow the experimental protocol in P3-SAM [ma2025p3]. For our method, a point prompt is provided to the model for each ground-truth part to generate its mask. These masks are then consolidated to yield a comprehensive segmentation for the entire object. Please refer to Suppl. for more information.
@@
Results. We report quantitative results of interactive segmentation in Table 1. Although our method can optionally incorporate scale prompts to boost performance, we also provide a non-scale version to ensure a fair comparison with the other two methods [zhou2024point, ma2025p3]. Table 1 shows that S2AM3D achieves the best results in both datasets [yang2024sampart3d, liu2023partslip]. With the input of scale prompts, our method further achieves a performance gain of 12.05% and 16.24%, demonstrating the effectiveness of our scale-aware decoding. The qualitative results in Figure 4 are basically consistent with the quantitative results. Our model produces segmentations largely consistent with the GT without scale condition; supplying scale guidance corrects the few cases of granularity mismatch by calibrating the level of detail.
@@
For full segmentation, Table 2 shows that S2AM3D achieves the best score on PartNet-E [liu2023partslip]. We also visualize the segmentation results in Figure 5. Constrained by 2D priors, methods like PartField [liu2025partfield] and SAMPart3D [yang2024sampart3d] produce segmentations with poor 3D consistency. On the other hand, 3D native methods like SAMesh [tang2024segment] and P3-SAM [ma2025p3] tend to fail in some long-tailed cases. In contrast, S2AM3D effectively circumvents these issues, achieving segmentation with higher completeness and better consistency on complex structures.

### v3

@@ Abstract
Part-level point cloud segmentation has recently attracted significant attention in 3D computer vision. Nevertheless, existing research is constrained by two major challenges: native 3D models lack generalization due to data scarcity, while introducing 2D pre-trained knowledge often leads to inconsistent segmentation results across different views. To address these challenges, we propose S2AM3D, which incorporates 2D segmentation priors with 3D consistent supervision. We design a point-consistent part encoder that aggregates multi-view 2D features through native 3D contrastive learning, producing globally consistent point features. A scale-aware prompt decoder is then proposed to enable real-time adjustment of segmentation granularity via continuous scale signals. Simultaneously, we introduce a large-scale, high-quality part-level point cloud dataset with more than 100k samples, providing ample supervision signals for model training. Extensive experiments demonstrate that S2AM3D achieves leading performance across multiple evaluation settings, exhibiting exceptional robustness and controllability when handling complex structures and parts with significant size variations. The project page is available at https://sumuru789.github.io/S2AM3D-website/.
@@ 1 Introduction
Part-level point cloud segmentation plays a pivotal role in bridging fine-grained geometric details with high-level semantic understanding, supporting significant applications in 3D content creation, robotic manipulation, and reverse engineering [37, 3, 9, 20, 17, 11], etc.Unlike instance detection, which only provides holistic semantics, part-level segmentation enables flexible granularity adjustment and seamless switching between local and global regions. This capability directly influences the feasibility and efficiency of downstream tasks, including part generation, replacement, assembly, and parametric editing. However, high-quality 3D part-level data remains scarce. Compared with the 2D domain, the cost of point cloud annotation is more prohibitive; consequently, existing datasets suffer from limited scale and category diversity, which severely hinders the generalization performance of 3D models [27, 28, 32, 15, 23].
@@
Recent works [19, 34, 18, 39, 40, 8, 30] leverage 2D pre-trained knowledge to enhance the generalization of 3D models, i.e., adopting 2D segmentation on 3D renderings. Segmented multi-view contents are then aggregated via mechanisms like multi-view lifting [8, 30, 19, 40] or distillation [34, 18, 39] to obtain the final 3D segmentation results. However, cross-view inconsistencies caused by occlusions, slender structures, and complex topology, can lead to accumulated errors that compromise global 3D consistency. To alleviate the inconsistency, some methods [18, 39] further incorporate 2D priors and provide supervision in the 3D space. Although native 3D data is used to enhance multi-view consistency, these methods still rely too heavily on 2D segmentation results. For example, PartField [18] primarily leverages part proposals from SAM [13], supplemented with limited 3D supervision to learn multi-view dependencies. Consequently, when 2D segmentation is affected by occlusions and other spatial factors, 3D representations struggle to correct these issues (see Figure 1).
@@
Extensive experiments are conducted to demonstrate the effectiveness of S2AM3D on part-level point cloud segmentation. As shown in Figure 1, our 2D-3D hybrid training recipe allows S2AM3D to achieve comparable performance to P3-SAM [23] with far less training data, and significantly outperforms SAMPart3D and PartField. Benefitting from the proposed scale-aware prompt decoder, S2AM3D exhibits remarkable flexibility by naturally extending its capability to point-prompted segmentation, making it a unified 3D part segmentation framework.
@@ 2 Related Work
Native 3D Part Segmentation. Early point cloud/mesh segmentation methods were typically trained under closed-label settings on ShapeNet-Part and PartNet, leveraging pointwise or neighborhood aggregation networks such as PointNet/PointNet++ and the Point Transformer family [27, 28, 38, 36, 25]. While strong within restricted categories and fixed granularities, these models often struggle under open-domain shapes, long-tailed parts, and cross-source shifts: (i) limited category/hierarchy coverage hampers generalization to unseen objects and rare parts; (ii) geometric-detail and noise discrepancies across sources obscure unified supervision; and (iii) the absence of explicit hierarchical consistency and granularity control constrains editing and interactive scenarios [2, 25, 4]. These limitations motivate scalable part understanding beyond closed-world assumptions.
@@
2D Transferred Part Segmentation. To mitigate the scarcity of native 3D annotations, many works transfer 2D foundation priors (e.g., SAM, CLIP/GLIP) to 3D [13, 29, 14]. Two mainstream paradigms emerge: (i) multi-view lifting renders shapes to images, applies 2D segmentation or vision–language priors per view, and fuses/back-projects/distills view-level cues into 3D supervision [13, 35, 33, 19]; and (ii) 2D-to-3D distillation directly trains native 3D networks from image-domain pseudo labels [10, 18, 34]. Despite reduced labeling cost and open-vocabulary semantics, both paradigms are prone to cross-view inconsistencies on thin/occluded structures, boundary ambiguity that undermines global 3D coherence, and costly post-processing pipelines; moreover, surface-only coverage limits supervision of internal structures [35, 33, 19, 10]. These observations point to the value of native 3D signals for globally consistent part supervision.
@@
Granularity Controllable Part Segmentation. Existing methods still face significant limitations in achieving flexible control over segmentation granularity. Feature clustering-based approaches (e.g., PartField [18] and SAMPart3D [34]) typically rely on post-processing clustering to determine segmentation granularity, resulting in non-continuous and non-intuitive control that hinders real-time fine-grained adjustment. Point prompt-based methods (e.g., Point-SAM [39] and P3-SAM [23]), while supporting interactive point prompting, lack explicit granularity control mechanisms and cannot precisely regulate segmentation scope through continuous parameters. Furthermore, the segmentation granularity in both types of methods remains fundamentally constrained by their reliance on 2D priors, making it challenging to ensure global consistency in 3D space. These limitations motivate us to explore a new solution that introduces an explicit continuous scale signal coupled with native 3D supervision, ultimately achieving truly reliable and controllable 3D segmentation granularity.
@@ 3.2 Point-Consistent Part Encoder
To aggregate point cloud features with general segmentation priors, the common practice is to train a point cloud encoder by distilling the pre-training knowledge from 2D segmentation models like SAM [13]. Following PartField  [18], we exploit a voxel-based encoder PVCNN [21] to extract the point latent and then convert it to the tri-plane [1] ($xy$, yzyz, and $zx$) representation with orthogonal projection. The Tri-plane features $\mathbf{T}\!\in\!\mathbb{R}^{3\times D\times H\times W}$ are then aggregated by a series of Transformer blocks, where $D$ is the feature channel. During training, tri-plane features are rendered from random viewpoints to formulate a 2D latent, which can be supervised with the distillation of SAM. However, relying solely on multi-view 2D distillation tends to introduce multi-view inconsistencies, leading to segmentation problems like boundary artifacts and cross-view conflicts.
@@
To obtain point features $\mathbf{F}$ given the 3D coordinate $(x,y,z)$ from the tri-plane representation [1], we unproject the point to the three feature planes of $\mathbf{T}$. The sampled features are summed across planes and stacked over all points as:
@@ 3.3 Scale-Aware Prompt Decoder
where $\{\omega_{k},\phi_{k}\}$ are learnable and $M$ denotes the number of sinusoidal frequency pairs. $\mathbf{e}(s)$ is then fed to a feature-wise linear modulation [26] (FiLM) in the channel dimension, which modulates the global feature map with a learnable gate $\alpha$:
$\left[\boldsymbol{\gamma},\boldsymbol{\beta}\right]$ $=\text{Linear}\!\big(\mathrm{LN}(\mathbf{e}(s))\big),$ (4)
@@
where Linear() is a linear layer applied after Layer Normalization on $\mathbf{e}(s)$, $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{D}$ are channel-wise FiLM parameters [26], and $\alpha\in\mathbb{R}$ is a learnable scalar gate. To maintain a consistent scale context across semantic levels, we interleave FiLM with Transformer blocks $T$:
@@ 3.4 Training Objective
We adopt a decoupled training scheme: our point-consistent part encoder is trained with the contrastive objective [12] in Sec. 3.2, while the scale-aware prompt decoder is optimized subsequently with the encoder frozen.
@@ 4.1 Scalable Data Pipeline
The data pipeline consists of three steps: part annotation, quality filtering, and connectivity refinement. Given a 3D object, we adopt surface-area–proportional sampling and assign the corresponding part label to every sampled point, which follows the part-annotation mining strategy of PartCrafter [16]. However, such an automatic strategy inevitably introduces erroneous annotation artifacts, which may mislead the model training. To filter out possible wrong labels, we further propose quality filtering and connectivity refinement.
In the quality filtering stage (see Figure 3(b)), we first manually curate a small-scale dataset containing examples of both reasonable and unreasonable part annotations. We then train a binary PointNet [27] validator on this dataset. Each point cloud is represented by $N$ points, each defined by a 4D vector $(x,y,z,l)$ encompassing normalized coordinates and its part-label index. Once trained, this validator is applied across the entire dataset to automatically screen and filter out samples with unreasonable annotations, ensuring that only high-quality labeled data proceed to subsequent stages.
We then perform connectivity refinement (see Figure 3(c)): artist annotations may group spatially disconnected yet semantically related regions under the same label, which degrades prompt-segmentation precision and hampers isolating local structures. For the points of each label, we compute the axis-aligned bounding-box diagonal length $d$ and set the DBSCAN [7] radius to $\epsilon=d\times\epsilon_{\text{factor}}$. If multiple spatial clusters are detected within a single label, we split it into distinct new labels. Finally, we filter and retain objects containing 2 to 50 parts to ensure dataset balance and suitability for part segmentation tasks.
@@ 4.2 Dataset Overview
We mainly collect 3D assets from Objaverse [5, 6], which is one of the largest 3D object datasets. We filter more than 100,000 point cloud instances from this dataset, spanning over 400 categories. This category advantage lies in covering open-world shapes from simple objects to complex machinery, supporting broader geometric and semantic diversity. In total, we annotate around 1.2 million part labels, with an average of 11 parts per object. As shown in Figure 3(a), the part count distribution spans 2-50, exhibiting a long-tailed pattern to ensure balanced coverage of fine- and coarse-grained shapes.
@@ 5.1 Implementation Details
Encoder training. The encoder is initialized with PartField [18] pre-trained parameters and then further optimized with contrastive learning on native point cloud annotations. For each mini-batch, we uniformly sample $\hat{N}=1000$ points to form $\hat{P}$ and compute the contrastive loss on it with temperature $\tau=0.07$. We use AdamW [22] with a learning rate of $1\times 10^{-5}$. The encoder is trained for 15 epochs, which takes about one day on 8 NVIDIA A6000 GPUs.
Decoder training. In this phase, we freeze the parameters of the encoder and fine-tune the decoder only. We set the layer number of scale modulation $L_{m}$ to 2 and the layer number of bi-directional cross-attention blocks $L_{d}$ to 4. The segmentation loss is a weighted sum of dynamic BCE and Dice, with $\lambda_{\mathrm{bce}}=0.7$ and $\lambda_{\mathrm{dice}}=0.3$. During training, we apply the scale dropout with a probability of 0.1, and at inference, we threshold the probability at $\theta=0.7$. We use AdamW [22] with a learning rate of $1\times 10^{-4}$. The training batch size is 40. The decoder is trained for 110 epochs with 4 NVIDIA A6000 GPUs.
@@ 5.2 Comparison
Point-SAM [39] 31.46 50.23 40.85
P3-SAM [23] 35.05 39.98 37.52
Ours 46.47 62.52 54.50
Ours (+scale) 61.19 77.51 69.35
@@
Figure 5: Qualitative comparison of full segmentation (PartObjaverse-Tiny [34]). For ease of comparison with our point cloud method, mesh-level outputs are presented as point clouds by uniformly sampling the segmented meshes.
@@
Find3D [24] 20.76 21.69 21.23
SAMPart3D [34] 48.79 56.17 52.48
SAMesh [31] - 26.66 -
PartField [18] 51.54 59.10 55.32
P3-SAM [23] 58.10 65.39 61.75
Ours 63.29 77.98 70.64
@@
Evaluation Datasets. We choose PartObjaverse-Tiny [34] and PartNet-[19] for experimental evaluation. PartObjaverse-Tiny [34] is a compact subset of Objaverse [6, 5], containing 200 samples across 8 categories with manual part annotations. PartNet-[19] provides 1,906 point cloud shapes across 45 categories, which can be used to assess cross-category generalization.
@@
Baselines. For interactive segmentation, we compare two point-prompt baselines: Point-SAM [39] and P3-SAM [23]. For fair comparison, all methods are evaluated with their default input configurations: Point-SAM [39] employs both coordinates and color information, while P3-SAM [23] utilizes coordinates and surface normals. In contrast, our approach requires only XYZ coordinates. For full segmentation, we select current state-of-the-art methods, including Find3D [24], SAMPart3D [34], SAMesh [31], PartField [18], and P3-SAM [23].
@@
Metrics. We use Intersection over Union (IoU) as the evaluation metric. The metric is computed across all parts of a single object to obtain the per-sample score, which is subsequently averaged over the entire dataset. For interactive segmentation, we follow the strategy in Point-SAM [39], which selects the prompt point for each part as the interior point farthest from its boundary. All methods are evaluated on 10,000 points normalized to the unit sphere. For full segmentation, we follow the experimental protocol in P3-SAM [23]. For our method, a point prompt is provided to the model for each ground-truth part to generate its mask. These masks are then consolidated to yield a comprehensive segmentation for the entire object. Please refer to Suppl. for more information.
@@
Results. We report quantitative results of interactive segmentation in Table 1. Although our method can optionally incorporate scale prompts to boost performance, we also provide a non-scale version to ensure a fair comparison with the other two methods [39, 23]. Table 1 shows that S2AM3D achieves the best results in both datasets [34, 19]. With the input of scale prompts, our method further achieves a performance gain of 14.72% and 14.99%, demonstrating the effectiveness of our scale-aware decoding. The qualitative results in Figure 4 are basically consistent with the quantitative results. Our model produces segmentations largely consistent with the GT without scale condition; supplying scale guidance corrects the few cases of granularity mismatch by calibrating the level of detail.
@@
For full segmentation, Table 2 shows that S2AM3D achieves 63.29 mIoU on PartObjaverse-Tiny and 77.98 mIoU on PartNet-[19]. We also visualize the segmentation results in Figure 5. Constrained by 2D priors, methods like PartField [18] and SAMPart3D [34] produce segmentations with poor 3D consistency. On the other hand, 3D native methods like SAMesh [31] and P3-SAM [23] tend to fail in some long-tailed cases. In contrast, S2AM3D effectively circumvents these issues, achieving segmentation with higher completeness and better consistency on complex structures.